OpenStudy (alex6799):

The table shows the estimated number of deer living in a forest over a five-year period. Which type of function best models the data? Write an equation to model the data. year Estimated Population 0 100 1 79 2 62 3 49 4 39

OpenStudy (alex6799):

<o:p></o:p> A. linear; y = 0.79x + 100 B. quadratic; y = 100x2 + 0.79 C. exponential; y = 100 * 0.79x D. quadratic; y = 0.79x2 + 100